# Towards Zero-Shot Multimodal Machine Translation

### v1

@@ Abstract
Current multimodal machine translation (MMT) systems rely on fully supervised data (i.e models are trained on sentences with their translations and accompanying images). However, this type of data is costly to collect, limiting the extension of MMT to other language pairs for which such data does not exist. In this work, we propose a method to bypass the need for fully supervised data to train MMT systems, using multimodal English data only. Our method, called ZeroMMT, consists in adapting a strong text-only machine translation (MT) model by training it on a mixture of two objectives: visually conditioned masked language modelling and the Kullback-Leibler divergence between the original and new MMT outputs. We evaluate on standard MMT benchmarks and the recently released CoMMuTE, a contrastive benchmark aiming to evaluate how well models use images to disambiguate English sentences. We obtain disambiguation performance close to state-of-the-art MMT models trained additionally on fully supervised examples. To prove that our method generalizes to languages with no fully supervised training data available, we extend the CoMMuTE evaluation dataset to three new languages: Arabic, Russian and Chinese. We further show that we can control the trade-off between disambiguation capabilities and translation fidelity at inference time using classifier-free guidance and without any additional data. Our code, data and trained models are publicly accessible. [^1] https://github.com/MatthieuFP/CoMMuTE [^2] https://github.com/MatthieuFP/zerommt
@@ 1 Introduction
Multimodal machine translation (MMT) refers to the use of additional modalities, such as images or videos, in machine translation (MT) systems. The main purpose is to provide an additional signal in the case of ambiguity in the text to be translated. Most current MMT models are trained solely on the Multi30K (M30K) dataset (Elliott et al., 2016, 2017; Barrault et al., 2018), a multilingual and multimodal corpus composed of 30,000 images, their captions in English and translations in French, German and Czech. There have been recent breakthroughs in MMT thanks to the use of pretrained text-only MT systems and monolingual captioning data in order to adapt MT systems to MMT (Futeral et al., 2023; Gupta et al., 2023; Vijayan et al., 2024). Good results have been shown using this strategy on CoMMuTE (Futeral et al., 2023), a contrastive benchmark designed to evaluate MMT models on their use of images to disambiguate between contrastive translations, and these results were significantly better than MMT systems trained on M30K only (Yin et al., 2020; Yao and Wan, 2020; Liu et al., 2021; Wu et al., 2021; Li et al., 2022b). However, these models still rely on the multilingual and multimodal M30K corpus during training to ensure good translation performance. This presents a fundamental limitation: collecting translations for captioning data is expensive [^3] Authors of M30K stated they spent €23,000 on the translation of the 30,000 English captions into German. and this restricts the extension of MMT to new languages. Attempts to bypass the problem have been made by carrying out zero-shot transfer between languages (Hirasawa et al., 2023), but this results in the poor use of images to disambiguate ambiguous texts (i.e. to correctly exploit the visual modality).
@@
In this work, we address this limitation by proposing a method requiring only monolingual multimodal text data, removing the need for fully supervised data, i.e. data such as M30K that is both parallel and multimodal. To do so, we start from a strong pretrained MT system and use it to translate multimodal English data into the target languages of interest. We then adapt the pretrained MT system to images using two objectives: (1) visually conditioned masked language modelling (VMLM) (Li et al., 2019; Lu et al., 2019) on multimodal English data to force the model to use image information and (2) a KL penalty on the translated multimodal data to maintain translation capabilities. We test our method on six languages directions: English to French, Czech, German, Arabic, Russian and Chinese, extending the CoMMuTE dataset to cover the three additional languages [^4] https://github.com/MatthieuFP/CoMMuTE. Our method, called ZeroMMT, obtains CoMMuTE scores close to the supervised state of the art, while there is only a small drop in BLEU and COMET scores compared to the underlying text-only MT system on standard MMT benchmarks composed mainly of unambiguous examples. We further show that we can control the trade-off between disambiguation and global translation performance at inference time with classifier-free guidance.
@@ 2 Related Work
Research in MMT originally focused on which visual features to use (Li et al., 2022a) and how to integrate them into sequence-to-sequence models (Sutskever et al., 2014) while training them from scratch on the widely used M30K benchmark (Libovický et al., 2016; Calixto et al., 2016; Elliott and Kádár, 2017; Calixto and Liu, 2017; Yao and Wan, 2020; Yin et al., 2020; Lin et al., 2020; Liu et al., 2021; Li et al., 2022b). These MMT systems typically show improvements of around 1-2 BLEU points on standard MMT benchmarks in comparison to text-only baselines that were also trained from scratch, which is not significant enough to state that MMT systems are better than their text-only counterparts (Mathur et al., 2020). Wu et al. (2021) additionally observed that while they obtained +1 BLEU on average on M30K test sets with the use of images, they got the same improvements with randomly initialized visual features, most likely due to regularization, i.e. the images were in reality not being exploited effectively. On top of that, being trained from scratch on fully supervised MMT data only, these models lag far behind state-of-the-art MT systems (Liu et al., 2020; Costa-jussà et al., 2022) trained on large amounts of parallel text data.
@@
Futeral et al. (2023) show that M30K contained few ambiguous examples that required visual context, and that MMT models can obtain decent translation performance on the benchmark while struggling to exploit images correctly. To handle this, they introduce VGAMT, an adapted MMT model building on top of a frozen state-of-the-art MT model. They also show that visually masked language modelling (VMLM) on English captioning data was a key additional objective to force MMT systems to become truly multimodal. Sato et al. (2023) and Bowen et al. (2024) further show that choosing the masked tokens in a smart way instead of randomly slightly boosts results. However, all these methods still require the use of fully supervised data to get good translation quality.
@@
There have been efforts to train MMT models without using fully supervised data (Su et al., 2019; Huang et al., 2020; Fei et al., 2023) in an unsupervised manner. These approaches are however fundamentally different from this work as their goal is to obtain MT models using synthetic text-only parallel data through the use of visual pivoting, not targeting disambiguation capabilities. Hirasawa et al. (2023) proposed a zero-shot method to learn MMT by training on the little fully supervised data available aiming for zero-shot cross-lingual transfer. As the amount of fully supervised data for a single language is small ($\leq$ 30K text-image pairs), and few languages are covered ($\leq$ 8), this method results in poor exploitation of the image to learn disambiguation capabilities.
@@
The main test sets to evaluate MMT systems are the test subsets of Multi30K (Elliott et al., 2016, 2017; Barrault et al., 2018). However, some of the translations were produced without access to the images and they have also been found to contain only a few ambiguous examples (Futeral et al., 2023) where visual context is necessary. They are therefore not best adapted to evaluating MMT systems when used alone. Elliott (2018) and Caglayan et al. (2019) further proposed to use an adversarial evaluation method and a probing method based on masked text inputs to assess the utility of images in the translation process. While this is a good sanity check, it is not a good proxy for evaluating the capacity of an MMT model to disambiguate translations. Lala and Specia (2018), Li et al. (2021) and Zhu et al. (2023) released evaluation datasets composed of sentences in English with ambiguous words accompanied with disambiguating images. However, Li et al. (2021) only target gender ambiguity, and all these datasets are prone to distributional bias, which is difficult to measure and such that text-only MT systems can perform very well on them. Traditional MT metrics (Papineni et al., 2002; Banerjee and Lavie, 2005; Rei et al., 2020) are also unable to catch how well MMT systems use images as they do not specifically target translations where images would be required to translate correctly. Tackling these issues, Futeral et al. (2023) introduced CoMMuTE, a contrastive evaluation dataset, composed of English sentences to be translated, built around ambiguous words, each accompanied by two translations with two images, each of which disambiguates the English sentence. They use perplexity to evaluate MMT models. Therefore text-only MT systems can only perform as well as random (50%), as they do not have access to the images, and it is also a direct proxy of MMT models’ ability to use the image to translate and disambiguate.
@@ 3 Extending the CoMMuTE benchmark
CoMMuTE (Futeral et al., 2023) is a contrastive test set designed to evaluate how well MMT systems use images to disambiguate ambiguous sentences and their translations. From an English sentence built around an ambiguous English word, two possible translations, each accompanied with one image are available. The associated image is chosen to disambiguate the English sentence such that one translation is correct and the other incorrect. Available for English-to-{French,German,Czech}, we extend CoMMuTE to three new target languages: Arabic, Chinese and Russian. [^5] Professional translators were asked to translate the English sentence for each of the two images. We also release a small validation set composed of 30 English ambiguous words [^6] Not overlapping with those from the test set. with two French translations, each associated with one image, to be used for model selection during training. Table 1 shows statistics of the extension of CoMMuTE. [^7] Token counts were calculated with the NLLB tokenizer (Costa-jussà et al., 2022).
@@ 4 Our Approach
Our goal is to train an MMT model capable of effectively using images to disambiguate contrasting translations while maintaining its translation capabilities, without using any fully supervised data. To do so, and as shown in Figure 1, we start from a strong pretrained MT system (different versions of NLLB (Costa-jussà et al., 2022) depending on the targeted model size), and use it to translate English captions into the target languages. Similarly to Futeral et al. (2023), we then turn it into an MMT model by adding lightweight trainable modules (visual projectors and adapters), keeping original weights frozen during training. We use SigLIP (Zhai et al., 2023) to provide visual embeddings that are concatenated to the sequences of text embeddings in the NLLB encoder. We then frame the problem of training an MMT model into two sub-objectives: (1) forcing the model to use images in the translation process using a visually-conditioned masked language modeling objective (VMLM) while (2) maintaining the performance of the original MT system without any fully supervised data using the Kullback-Leibler (KL) divergence between the MMT system’s and the original MT system’s output distributions using the previously automatically translated data. We call our approach ZeroMMT.
Concretely, let $x_{1,\dots,n}$ denote the sequence of tokens of the English sentence, $i$ the image embedding, $y_{1,\dots,m}$ the translated sequence of tokens, $f_{\theta}$ the original MT system and $f_{\theta,\beta}$ the MMT system built on top of the text-only MT model with additional light-weight modules $\beta$, both outputting probability distributions over tokens. We formally define the losses as follows: $\mathscr{L}_{\textit{VMLM}}=\sum_{j}y_{j}\log\big{(}f_{\theta,\beta}(y_{j};y_{<j},x_{\setminus\mathscr{M}},i)\big{)}$ (1) $\mathscr{L}_{\textit{KL}}=\sum_{j}f_{\theta}(y_{j};y_{<j},x)\log\frac{f_{\theta}(y_{j};y_{<j},x)}{f_{\theta,\beta}(y_{j};y_{<j},x,i)}$ (2)
where $\mathscr{M}$ is the set of masked input indices. The final loss is a weighted combination of (1) and (2), and we choose $\lambda$ value based on results on validation sets as described in Section 5.2:
@@ 5.1 Data
We trained our models on the Conceptual Captions dataset [^8] At the time of writing, we were able to collect 2,831,746 out of the 3,300,000 images. (Sharma et al., 2018). We translated Conceptual Captions into French, German, Czech, Chinese, Russian and Arabic using the underlying NLLB (Costa-jussà et al., 2022) MT system (of size 600M, 1.3B or 3.3B depending on the experiment) using beam search (Graves, 2012) with a beam of size 4 for the 600M model and 2 for the largest ones.
We evaluated our models on the Multi30K test sets (Elliott et al., 2016, 2017; Barrault et al., 2018) for English-to-{German,French,Czech}, the EMMT test set (Zhu et al., 2023) for English-to-Chinese, comprising 500 titles of commercial products from e-commercial websites translated from English to Chinese, the VATEX test set (Wang et al., 2019) for English-to-Chinese which is composed of 10-second videos [^9] We take 5 frames per second, compute SIGLIP features and average them to obtain the visual input. with English captions translated into Chinese, and finally CoMMuTE (Futeral et al., 2023) for English-to-{German,French,Czech,Chinese,Russian,Arabic}, used to test how well the MMT models exploit visual context for disambiguation.
@@
Ar Cs De Fr Ru Zh
Text-only MT baselines 50.0 50.0 50.0 50.0 50.0 50.0
NLLB-SIGLIP topline 82.6 76.0 83.7 75.0 75.8 88.1
MMT – fully supervised
Gated Fusion bilingual - 51.0 $\pm$1.9 49.7 $\pm$0.6 50.0 $\pm$0.8 - -
VTLM + MMT bilingual - 52.0 $\pm$0.7 50.2 $\pm$0.3 51.4 $\pm$0.9 - -
VGAMT full bilingual - 55.6 $\pm$0.8 59.0 $\pm$0.5 67.1 $\pm$0.7 - -
VGAMT SIGLIP-only multi. - 57.5 $\pm$1.2 57.1 $\pm$0.4 61.3 $\pm$1.1 - -
MMT – zero-shot
Table 3: Results on CoMMuTE, averaged over 3 runs ($\pm$ standard error).

### v2

@@ Abstract
Current multimodal machine translation (MMT) systems rely on fully supervised data (i.e sentences with their translations and accompanying images), which is costly to collect and prevents the extension of MMT to language pairs with no such data. We propose a method to bypass the need for fully supervised data to train MMT systems, using multimodal English data only. Our method (ZeroMMT) consists in adapting a strong text-only machine translation (MT) model by training it jointly on two objectives: visually conditioned masked language modelling and the Kullback-Leibler divergence between the original MT and new MMT outputs. We evaluate on standard MMT benchmarks and on CoMMuTE, a contrastive test set designed to evaluate how well models use images to disambiguate translations. ZeroMMT obtains disambiguation results close to state-of-the-art MMT models trained on fully supervised examples. To prove that ZeroMMT generalizes to languages with no fully supervised training data, we extend CoMMuTE to three new languages: Arabic, Russian and Chinese. We also show that we can control the trade-off between disambiguation capabilities and translation fidelity at inference time using classifier-free guidance and without any additional data. Our code, data and trained models are publicly accessible. [^1] https://github.com/MatthieuFP/CoMMuTE, [^2] https://github.com/MatthieuFP/zerommt
Towards Zero-Shot Multimodal Machine Translation
Matthieu Futeral1,2   Cordelia Schmid1,2   Benoît Sagot1   Rachel Bawden1 1Inria, Paris 2Département d’informatique de l’ENS, CNRS, PSL Research University firstname.lastname@inria.fr
@@ 1 Introduction
Multimodal machine translation (MMT) refers to the use of additional modalities, such as images or videos, in machine translation (MT) systems. The main purpose is to provide an additional signal in the case of ambiguity in the text to be translated (i.e the text alone does not provide enough information). Most current MMT models are trained solely on the Multi30K (M30K) dataset (Elliott et al., 2016, 2017; Barrault et al., 2018), a multilingual and multimodal corpus composed of 30K images, their English captions and translations in French, German and Czech. There have been recent breakthroughs in MMT thanks to the use of pretrained text-only MT systems and monolingual captioning data in order to adapt MT systems to MMT (Futeral et al., 2023; Gupta et al., 2023; Vijayan et al., 2024). Good results have been shown using this strategy on CoMMuTE (Futeral et al., 2023), a benchmark designed to evaluate MMT models on their use of images to disambiguate between contrastive translations, and these results were significantly better than MMT systems trained on M30K only (Yin et al., 2020; Yao and Wan, 2020; Liu et al., 2021; Wu et al., 2021; Li et al., 2022b). However, these models still rely on the multilingual and multimodal M30K corpus during training to ensure good translation performance. This presents a core limitation: collecting translations of captions is costly, [^3] Authors of M30K stated they spent €23,000 on the translation of the 30,000 English captions into German. restricting MMT’s extension to new languages. Zero-shot transfer between languages has been tested to bypass the problem (Hirasawa et al., 2023), but this results in the poor exploitation of the visual modality to disambiguate ambiguous texts.
@@
In this work, we address this limitation by proposing a method requiring only monolingual multimodal text data (i.e. English text-image pairs), removing the need for fully supervised data, i.e. parallel and multimodal data such as M30K. We start from a strong pretrained MT system and use it to translate multimodal English data into the target languages of interest. We then adapt the pretrained MT system to images using two objectives: (1) visually conditioned masked language modelling (VMLM) (Li et al., 2019; Lu et al., 2019) on multimodal English data to force the model to use image information and (2) a KL penalty on the translated multimodal data to maintain translation capabilities. We test our method on six languages directions: English to French, Czech, German, Arabic, Russian and Chinese, extending the CoMMuTE dataset to cover the three additional languages. Our method, called ZeroMMT, obtains CoMMuTE scores close to the supervised state of the art, while there is only a small drop in BLEU and COMET scores compared to the underlying text-only MT system on standard MMT benchmarks composed mainly of unambiguous examples (i.e. where images are not useful for correct translation). We further show that we can control the trade-off between disambiguation and general translation performance at inference time with classifier-free guidance.
@@ 2 Related Work
Research in MMT originally focused on which visual features to use (Li et al., 2022a) and how to integrate them into sequence-to-sequence models (Sutskever et al., 2014) trained from scratch on the widely used M30K benchmark (Libovický et al., 2016; Calixto et al., 2016; Elliott and Kádár, 2017; Calixto and Liu, 2017; Yin et al., 2020; Liu et al., 2021; Li et al., 2022b). These MMT systems typically show improvements of around 1-2 BLEU points on standard MMT benchmarks in comparison to text-only baselines trained from scratch, which is not significant enough to state that MMT systems are better than their text-only counterparts (Mathur et al., 2020). Wu et al. (2021) observed that while they obtained +1 BLEU on average on M30K test sets with the use of images, they got the same improvements with randomly initialized visual features, most likely due to regularization, i.e. the images were in reality not being exploited effectively. On top of that, being trained from scratch on fully supervised MMT data only, these models lag far behind state-of-the-art MT systems (Costa-jussà et al., 2022) trained on large amounts of parallel text.
@@
Futeral et al. (2023) show that M30K contains few ambiguous examples requiring visual context, and that models can get good results on the benchmark while still struggling to exploit images correctly. They introduce VGAMT, an adapted MMT model based on a frozen state-of-the-art MT model. They also show that visually masked language modelling (VMLM) on English captioning data was a key additional objective to force MMT systems to become truly multimodal. Sato et al. (2023) and Bowen et al. (2024) further show that choosing the masked tokens in a smart way instead of randomly slightly boosts results. However, these methods still require fully supervised data to be good at translation; training on VMLM alone results in a collapse in translation capabilities.
A few works have used pseudo-multimodal parallel data by translating English captions into the target language using a pretrained MT system (Li et al., 2021; Caglayan et al., 2021; Vijayan et al., 2024). However, Caglayan et al. (2021) and Vijayan et al. (2024) used them in a pretraining step in a form of distillation of the knowledge of the MT system into the new MMT model before fine-tuning on M30K. Li et al. (2021) use backtranslation to translate English captions into Turkish to train a Turkish-to-English MMT model for disambiguating gender pronouns from Turkish to English. While effective, their method cannot be applied beyond this particular context because it requires the MT system to output the correct translation, which cannot be assumed to be the case in more general ambiguous contexts (i.e. when text context is not enough to translate the English text correctly).
@@
There have been efforts to train MMT models without using fully supervised data (Su et al., 2019; Huang et al., 2020; Fei et al., 2023). These approaches are however fundamentally different from this work as their goal is to obtain MT models using synthetic text-only parallel data through the use of visual pivoting, not targeting disambiguation capabilities. Hirasawa et al. (2023) proposed a zero-shot method to learn MMT by training on the little fully supervised data available aiming for zero-shot cross-lingual transfer. As the amount of fully supervised data for a single language is small ($\leq$ 30K text-image pairs), and few languages are covered ($\leq$ 8), this method results in poor exploitation of the image to learn disambiguation capabilities.
@@
The test sets typically used to evaluate MMT systems are the test subsets of M30K (Elliott et al., 2016, 2017; Barrault et al., 2018). However, some of the translations were produced without access to the images and they have also been found to contain only a few ambiguous examples where visual context is necessary (Futeral et al., 2023). They are therefore not best adapted to evaluating MMT systems. Elliott (2018) and Caglayan et al. (2019) proposed to use an adversarial evaluation method and a probing method based on masked text inputs to assess the utility of images in translation. However, this is not a good proxy for evaluating an MMT model’s capacity to disambiguate translations as it relies on masked inputs and so says little about models’ capacities to disambiguate when given unmasked text inputs. Lala and Specia (2018), Li et al. (2021) and Zhu et al. (2023) released evaluation datasets composed of sentences in English with ambiguous words accompanied with disambiguating images. However, Li et al. (2021) only target the ambiguity of gender pronouns, and all these datasets are prone to distributional bias, which is difficult to measure and is such that text-only MT systems can perform very well on them (i.e. images are in fact often not necessary for correct translation). Traditional MT metrics (Papineni et al., 2002; Banerjee and Lavie, 2005; Rei et al., 2020) are also unable to catch how well MMT systems use images as they do not specifically target translations where images would be required to translate correctly. Tackling these issues, Futeral et al. (2023) introduced CoMMuTE, a contrastive evaluation dataset, composed of English sentences to be translated, built around ambiguous words, each sentence accompanied by two translations with two images, each of which disambiguates the English sentence. MMT models are evaluated on their capacity to give a lower perplexity to the correct translation than the incorrect one, given the source sentence and an image. As perplexity is used to evaluate MMT models, it is a direct proxy of MMT models’ capacity to disambiguate English sentences. Furthermore, text-only MT systems can only perform as well as random (50%), as they do not have access to the images.
@@ 3 Extending the CoMMuTE benchmark
Currently available for English-to-{French,German,Czech}, we extend the CoMMuTE benchmark to three new target languages: Arabic, Chinese and Russian, using professional translators. We also release a small validation set of 30 English ambiguous words (non-overlapping with test set examples) with two French translations, each with its own image, to be used for model selection during training. Table 1 shows statistics of our extension of CoMMuTE. [^4] Tokenisation using NLLB (Costa-jussà et al., 2022).
@@ 4 Our Approach
Our goal is to train an MMT model capable of effectively using images to disambiguate ambiguous translations (i.e. where an image is necessary to translate correctly, which is MMT’s main purpose) while keeping the general MT capacity of the underlying MT model, without using fully supervised data (i.e. in a zero-shot way). This allows us to extend MMT to more language pairs, currently not possible without collecting fully supervised data.
As shown in Figure 1, we start from a strong pretrained NLLB (Costa-jussà et al., 2022) MT model and use it to translate English captions. Similarly to Futeral et al. (2023), we turn it into an MMT model by adding lightweight trainable modules (visual projectors and adapters), keeping original weights frozen during training. We use visual embeddings from SigLIP (Zhai et al., 2023) and concatenate them to the sequences of text embeddings in the NLLB encoder. Our approach, ZeroMMT, is based on two objectives: (1) force the model to use images when translating by using visually-conditioned masked language modeling (VMLM) and (2) maintain the performance of the original MT system without any fully supervised data using the Kullback-Leibler (KL) divergence between the MMT system’s output and the original MT system’s output distributions using the previously automatically translated data. While (1) has already been proved successful for learning visual disambiguation capabilities (Futeral et al., 2023), we further show (Section 6) that (2) is key for retaining a strong translation capacity in an MMT setting, enabling zero-shot training.
In more detail, let $x_{1,\dots,n}$ denote the sequence of tokens of the English sentence, $i$ the image embedding, $y_{1,\dots,m}$ the translated sequence of tokens, $f_{\theta}$ the original MT system and $f_{\theta,\beta}$ the MMT system built on top of the text-only MT model with additional light-weight modules $\beta$, both outputting probability distributions over tokens. We formally define the losses as follows: $\mathscr{L}_{\textit{VMLM}}=\sum_{j}y_{j}\log\big{(}f_{\theta,\beta}(y_{j};y_{<j},x_{\setminus\mathscr{M}},i)\big{)}$ (1) $\mathscr{L}_{\textit{KL}}=\sum_{j}f_{\theta}(y_{j};y_{<j},x)\log\frac{f_{\theta}(y_{j};y_{<j},x)}{f_{\theta,\beta}(y_{j};y_{<j},x,i)}$ (2)
where $\mathscr{M}$ is the set of masked input indices. The final loss is a weighted combination of (1) and (2), and we choose the $\lambda$ value based on results on validation sets as described in Section 5.2:
@@ 5.1 Data
We trained our models on the Conceptual Captions dataset [^5] At the time of writing, we were able to collect 2,831,746 out of the 3,300,000 images. (Sharma et al., 2018). We translated Conceptual Captions into French, German, Czech, Chinese, Russian and Arabic using NLLB (Costa-jussà et al., 2022) (of size 600M, 1.3B or 3.3B depending on the experiment) using a beam of size 4 for the 600M model and 2 for the largest ones.
We evaluate our models on the M30K test sets (Elliott et al., 2016, 2017; Barrault et al., 2018) for English-to-{German,French,Czech}, the EMMT test set (Zhu et al., 2023) for English-to-Chinese, comprising 500 English product titles from e-commercial websites translated into Chinese, and the VATEX test set (Wang et al., 2019) for English-to-Chinese, composed of 10-second videos [^6] We take 5 frames per second, compute SIGLIP features and average them to obtain the visual input. with English captions translated into Chinese. We use these test sets to make sure general translation quality is not harmed when introducing additional visual inputs in unambiguous cases (as described previously, they cannot be used in practice to evaluate MMT models’ ability to use images correctly). Finally we evaluate on CoMMuTE for English-to-{German,French,Czech,Chinese,Russian,Arabic}, used to test how well the MMT models exploit visual context for disambiguation.
Ar Cs De Fr Ru Zh
Text-only MT baselines 50.0 50.0 50.0 50.0 50.0 50.0
NLLB-SIGLIP topline 82.6 76.0 83.7 75.0 75.8 88.1
MMT – fully supervised
Gated Fusion bilingual - 51.0 $\pm$1.9 49.7 $\pm$0.6 50.0 $\pm$0.8 - -
VTLM + MMT bilingual - 52.0 $\pm$0.7 50.2 $\pm$0.3 51.4 $\pm$0.9 - -
VGAMT full bilingual - 55.6 $\pm$0.8 59.0 $\pm$0.5 67.1 $\pm$0.7 - -
VGAMT SIGLIP-only multi. - 57.5 $\pm$1.2 57.1 $\pm$0.4 61.3 $\pm$1.1 - -
MMT – cross-lingual zero-shot
M2KT-VPN bilingual - 50.1 $\pm$0.6 50.3 $\pm$0.5 50.9 $\pm$0.8 - -
MMT – zero-shot
Table 2: Results on CoMMuTE, averaged over 3 runs ($\pm$ standard error). The best scores for each category are in bold and the second best are underlined.